OpenStudy (anonymous):

Solve 3^x – 4 = 7^x + 9

OpenStudy (ash2326):

@trumpetmaster7777 Add 4 to both the sides

OpenStudy (anonymous):

so it would be wait wat? the x-4 is to the power.

OpenStudy (anonymous):

They're exponential equations.

OpenStudy (ash2326):

is this your question? \[3^x -4=7^x+9\] or \[3^{x-4}=7^{x+9}\]

OpenStudy (anonymous):

the second one

OpenStudy (ash2326):

@trumpetmaster7777 Any power of 3 will be a multiple of 3 which can never have a factor of 7. Do you agree?

OpenStudy (anonymous):

i guess.

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

right, so x can't be an integer. Let's use logarithm here. Do you know logarithm?

OpenStudy (anonymous):

no

OpenStudy (ash2326):

Have you studied it ever? without it I'm not sure how can I help

OpenStudy (anonymous):

crap, guess i'll just guess.

OpenStudy (anonymous):

get it?

OpenStudy (ash2326):

let me think

OpenStudy (anonymous):

k

OpenStudy (ash2326):

Could you check the question again?